# Linear-Size Ancilla Systems for Logical Measurements in QLDPC Codes

### v1

@@ 3.6 Gauged $Y$ Ancilla System
If the $X$-system and $Z$-system individually have code distances at least $d$, then the $Y$-system has code distance at least $d$.
Proof.
Suppose $\bar{P}$ is a nontrivial logical operator of $\mathcal{G}_{Y}$, the merged code of the $Y$ system. We aim to show that it has weight at least $d$. Let $\bar{P}=P_{X}P_{Z}$ be its decomposition into an $X$-type Pauli $P_{X}$ and $Z$-type Pauli $P_{Z}$.
Throughout we let $\mathcal{G}_{X}$ and $\mathcal{G}_{Z}$ denote the merged codes of the $X$- and $Z$-systems, respectively. When we write $P^{\prime}_{X}=P_{X}|_{\mathcal{G}_{X}}$ or $P^{\prime}_{Z}=P_{Z}|_{\mathcal{G}_{Z}}$, we mean to restrict those operators originally supported on $\mathcal{G}_{Y}$ to the qubits that exist in $\mathcal{G}_{X}$ or $\mathcal{G}_{Z}$, respectively.
We note that $P^{\prime}_{X}$ is a logical operator on $\mathcal{G}_{X}$. It commutes with all $Z$ checks of $\mathcal{G}_{X}$ because $\bar{P}$ commutes with all the checks of $\mathcal{G}_{Y}$. Similarly, $P^{\prime}_{Z}$ is a logical operator of $\mathcal{G}_{Z}$.
We next claim that either $P^{\prime}_{X}$ is a nontrivial logical operator of $\mathcal{G}_{X}$ or $P^{\prime}_{Z}$ is a nontrivial logical operator of $\mathcal{G}_{Z}$. This can be seen by contradiction. Assuming both $P^{\prime}_{X}$ and $P^{\prime}_{Z}$ are products of stabilizers in their respective codes, then $\bar{P}$ can also be multiplied by the respective stabilizers from $\mathcal{G}_{Y}$ to find an equivalent operator with support that is much restricted:
where we used $C^{Z}_{j\text{ odd}}$ to mean $\bigcup_{j\text{ odd}}C^{Z}_{j}$ and similarly for the other variants of this notation. Furthermore, by multiplying with gauge checks $U^{X},U^{Z},U^{B}$ and checks from $C^{Z}_{j\text{ odd}}$ and $C^{X}_{i\text{ even}}$ we can move the support of $\bar{P}\equiv_{\mathcal{S}_{Y}}Z(b^{\prime}_{Z}\in B)X(b^{\prime}_{X}\in B)$ entirely onto the bridge qubits $B$ (this goes like the proof of Theorem 3 so we omit the details). However, in order to commute with all checks in $U^{B}$ (see Lemma 16), $V_{1}^{Z}$, and $V_{1}^{X}$, it must be that $b^{\prime}_{Z}=b^{\prime}_{X}=0$ or $\bar{P}\equiv_{\mathcal{S}_{Y}}I$. This is impossible because then $\bar{P}$ would be a nontrivial logical operator of $\mathcal{G}_{Y}$.
Now to conclude the proof, we simply note that $P^{\prime}_{X}$ being nontrivial in $\mathcal{G}_{X}$ implies $|P^{\prime}_{X}|\geq d$ by assumption. Since $|\bar{P}|\geq|P_{X}|\geq|P^{\prime}_{X}|\geq d$, we obtain the desired lower bound on $|\bar{P}|$ in this case. Likewise, if $P^{\prime}_{Z}$ is nontrivial in $\mathcal{G}_{Z}$, we find $|\bar{P}|\geq|P_{Z}|\geq|P^{\prime}_{Z}|\geq d$. ∎
@@ 4 Case Study: Gross Code
Figure 8: Qubit and degree counts of a gross code equipped with an ancilla system. This ancilla system implements eight different logical measurements by realising different subgraphs as Tanner graphs. Circular qubits are data, and squares are checks. Gauge checks are in purple, the bridge is green, and the remaining ancilla system qubits are colored according to their partners in the gross code: X is red, Z is teal, R is oRange, and L is bLue. The ancilla system is connected to two pairs of logical operators that are ZX-duals of one another: the $\bar{X},\bar{Z}$ qubits are opaque and the $\bar{X}^{\prime},\bar{Z}^{\prime}$ qubits are transparent. Their Tanner-subgraphs overlap on the $ZZ^{\prime}_{0}$ and $XX^{\prime}_{0}$ qubits. Note that checks are labeled $X$ or $Z$ type according to how they are used in $X$ or $Z$ systems individually, but in the joint measurement and $Y$ systems that employ the bridge, they are no longer mono-type checks (see Figs. 5 and 7). An annotation ‘3 d7’ means that there are three degree-seven qubits.
@@ 4.1 Mono-layer ancilla system
Several gauge checks were introduced, either supported in the $X$ or $Z$ systems alone, or to deal with loops introduced by the bridge. The gauge checks are specified in Figures 9 and 10. For the full $\bar{X}\bar{X}^{\prime}$ and $\bar{Z}\bar{Z}^{\prime}$ measurement operations all of these gauge checks are necessary. But for the other measurements, it turns out that several of the gauge checks are redundant. In the $\bar{X},$ and $\bar{Z}^{\prime}$ systems one only needs gauge checks 0, 1, and 2. In the $\bar{Z}$ and $\bar{X}^{\prime}$ systems one only needs gauge check $6$. In the $\bar{Y}$ and $\bar{Y}^{\prime}$ systems, gauge checks 10 and 17 must have their support from bridge qubit 0 removed in order to be well-defined. Then, gauge checks 6, 12, 16, 17, 18, and 20 are redundant.
@@ Acknowledgements
We are grateful to Ben Brown, John Blue, and Michael Beverland for enlightening discussions and valuable feedback.

### v3

@@ 3.8 Bridging Codes
Remark 25.
We say that this is a version of our fault-tolerance guarantee because the assumption on $\bar{X}_{1},\bar{X}_{2}$ needed for this theorem to hold is flexible. In the current statement, we ask that $\bar{X}_{1}$ and $\bar{X}_{2}$ are minimal weight representatives to account for edge cases. If we assume that their induced Tanner graphs are expanding and their ancilla systems have enough layers, then fault distance of the joint system can be proved as in Theorem 17.
This result enables us to connect different families of QLDPC codes together for a fault-tolerance architecture. Similar ideas based on the CKBB scheme have been considered and studied before [CKBB22, BCG+24, XBAP+24]. However, these constructions mostly focused on connecting a QLDPC code with a surface code, and lose fault-tolerant when we use fewer than $O(d)$ layers in the ancilla system.
Our bridge construction overcomes this barrier and can be used as a code adapter: two arbitrary code blocks can be fault-tolerantly connected, enabling teleportation of logical information. This capability opens many possibilities for a QLDPC-based computing architecture. In particular, we can connect a QLDPC code with efficient logical Clifford operators to a magic state factory based on a vastly different code, achieving universality by combining individually optimized methods.
On the other hand, our bridge construction comes with a noteworthy caveat. As discussed in Section 3.6, the weight of the gauge checks $U^{B}$ depends on the logical operators being measured, which means in the worst case these checks may have high weight. Let us discuss the choice of $U^{B}$ in more detail. Let $R$ be the repetition code on $B$ bits, from Lemma 15, we see that the checks $U^{B}$ restricted to the bridge qubits generate the checks of a repetition code, namely $R^{\perp}$. Given the induced Tanner graph of a logical operator, we can choose the connecting matrices $S^{(1)},S^{(2)}$ and a basis for $R^{\perp}$ to minimize the weight of $U^{B}$. This is captured in the following formulation.
Problem 26.
Find a low weight $(|B|-1)\times|B|$ matrix $C_{R}$ with $\text{rowspace}(C_{R})=R^{\perp}$, such that there exist low weight matrices $T^{(1)},T^{(2)}$ and connecting matrices $S^{(1)},S^{(2)}$ satisfying $T^{(1)}F^{(1)}=C_{R}S^{(1)^{\top}}$ and $C_{R}S^{(2)^{\top}}=T^{(2)}F^{(2)}$.
In the case where $F^{(1)}=F^{(2)}$, we can simply take $C_{R}$ to be a submatrix of $F^{(1)}$. In average cases, we expect reasonable choices of $C_{R},S^{(1)},S^{(2)}$ to exist. For arbitrary $F^{(1)}$ and $F^{(2)}$, however, such choices may not exist, and our bridge cannot connect these logical operators without breaking the LDPC property.
This limitation is fully resolved in [SJOY24], where it was shown that when $F^{(1)}$, $F^{(2)}$ are simple expander graphs (such as those constructed in [WY24, IGND24]), one can always efficiently solve Problem 26 to guarantee a LDPC bridge construction. This means the adapters constructed in [SJOY24] are truly universal, in the sense that they can connect arbitrary QLDPC code families while maintaining fault-tolerance and LDPC. We refer readers to [SJOY24] for further studies and applications of the idea of code adapting.
@@ 4.1 Mono-layer ancilla system
Figure 7: Qubit and degree counts of a gross code equipped with an ancilla system. This ancilla system implements eight different logical measurements by realising different subgraphs as Tanner graphs. Circular qubits are data, and squares are checks. Gauge checks are in purple, the bridge is green, and the remaining ancilla system qubits are colored according to their partners in the gross code: X is red, Z is teal, R is oRange, and L is bLue. The ancilla system is connected to two pairs of logical operators that are ZX-duals of one another: the $\bar{X},\bar{Z}$ qubits are opaque and the $\bar{X}^{\prime},\bar{Z}^{\prime}$ qubits are transparent. Their Tanner-subgraphs overlap on the $ZZ^{\prime}_{0}$ and $XX^{\prime}_{0}$ qubits. Note that checks are labeled $X$ or $Z$ type according to how they are used in $X$ or $Z$ systems individually, but in the joint measurement and $Y$ systems that employ the bridge, they are no longer mono-type checks (see Figs. 4 and 5). An annotation ‘3 d7’ means that there are three degree-seven qubits.
@@
Several gauge checks were introduced, either supported in the $X$ or $Z$ systems alone, or to deal with loops introduced by the bridge. The gauge checks are specified in Figures 8 and 9. For the full $\bar{X}\bar{X}^{\prime}$ and $\bar{Z}\bar{Z}^{\prime}$ measurement operations all of these gauge checks are necessary. But for the other measurements, it turns out that several of the gauge checks are redundant. In the $\bar{X},$ and $\bar{Z}^{\prime}$ systems one only needs gauge checks 0, 1, and 2. In the $\bar{Z}$ and $\bar{X}^{\prime}$ systems one only needs gauge check $6$. In the $\bar{Y}$ and $\bar{Y}^{\prime}$ systems, gauge checks 10 and 17 must have their support from bridge qubit 0 removed in order to be well-defined. Then, gauge checks 6, 12, 16, 17, 18, and 20 are redundant.
@@ Acknowledgements
We are grateful to Ben Brown, John Blue, Michael Beverland, Christopher Pattison and Dominic Williamson for enlightening discussions and valuable feedback. Z. He is supported by National Science Foundation Graduate Research Fellowship under Grant No. 2141064. This work was done in part while Z. He was visiting the Simons Institute for the Theory of Computing, supported by NSF QLCI Grant No. 2016245.
@@ A.2 Proofs for the $Y$ Ancilla System
See 22
Proof.
We will show that $d(\bar{X}_{M}\mathcal{L}\mathcal{U})\geq d$. The proof for $d(\bar{Z}_{M}\mathcal{L}\mathcal{U})\geq d$ follows similarly. As before, let $\mathcal{G}_{X}$ and $\mathcal{G}_{Z}$ denote the merged codes of the $X$- and $Z$-systems, respectively. For an operator $\bar{P}$, let $P_{X},P_{Z}$ denote its $X$ and $Z$ components, and let $P_{X}^{\prime}=P_{X}|_{\mathcal{G}_{X}},P_{Z}^{\prime}=P_{Z}|_{\mathcal{G}_{Z}\cup B}$. Due to the generators of $\operatorname{Init}$, we have
$d(\bar{X}_{M}\mathcal{L}\mathcal{U})=\min_{\bar{P}\in\bar{X}_{M}\mathcal{L}\mathcal{S}_{Y}}\mathcal{R}_{\operatorname{Init}}(\bar{P})=\min_{\bar{P}\in\bar{X}_{M}\mathcal{L}\mathcal{S}_{Y}}|P_{X}^{\prime}|+|P_{Z}^{\prime}|.$ (87)
Denote by $q_{1}^{X}$ and $q_{1}^{Z}$ the $X$ and $Z$ parts of the check $q_{1}=q_{1}^{X}q_{1}^{Z}$, we see that $\bar{X}_{M}\equiv_{\mathcal{S}_{Y}}Z(B)q_{1}^{Z}$. Consider stabilizers in the ancilla system, we observe that all gauge checks and all checks on even layers does not affect $|P_{X}^{\prime}|$ or $|P_{Z}^{\prime}|$. Therefore, let $\tilde{S}=\langle\mathcal{S}_{O},V^{X}_{i\text{ odd}},V^{Z}_{i\text{ odd}},q_{1}\rangle$, we have
$\min_{\bar{P}\in\bar{X}_{M}\mathcal{L}\mathcal{S}_{Y}}|P_{X}^{\prime}|+|P_{Z}^{\prime}|=\mathcal{R}_{\mathcal{L}\tilde{S}}(Z(B)q_{1}^{Z}).$ (88)
It suffices for us to show that the right hand side is at least $d$. Suppose we multiply $Z(B)q_{1}^{Z}$ by a collection of checks $v_{i}\in V_{i}^{X}$, $u_{j}\in V_{j}^{Z}$ for $i,j$ odd, and $(q_{1})^{s}$ for $s=0$ or $1$. Denote this new operator as $P$, namely,
$P=Z(B)q_{1}^{Z}\prod_{\begin{subarray}{c}i=1\\ i\text{ odd}\end{subarray}}^{L_{X}}\mathcal{H}(v_{i})\cdot\prod_{\begin{subarray}{c}j=1\\ j\text{ odd}\end{subarray}}^{L_{Z}}\mathcal{H}(u_{j})\cdot(q_{1})^{s}.$ (89)
To simplify notation, we collect the terms $q_{1}^{Z}$ and $(q_{1})^{s}$ in the following way: if $s=0$, we add vertex $q_{1}$ to $u_{1}$ (because effectively $P$ has a factor of $q_{1}^{Z}$); if $s=1$, we add vertex $q_{1}$ to $v_{1}$ (because effectively $P$ has a factor of $q_{1}^{X}$). Then by the Expansion Lemma
$|P|$ $\geq|B+S^{X}v_{1}+S^{Z}u_{1}|+|v_{m}+v_{L_{X}}|+|u_{m}+u_{L_{Z}}|+|v_{1}+v_{m}|+|u_{1}+u_{m}|$
$+\min(|v_{m}|,|\bar{X}|-|v_{m}|)+\min(|u_{m}|,|\bar{Z}|-|u_{m}|)+|v_{1}|+|u_{1}|.$ (90)
$P$ now has $X$-support $v_{1}\in V_{0}^{X}$ and $Z$-support $u_{1}\in V_{0}^{Z}$. If we further multiply $P$ by stabilizers in $\mathcal{S}_{O}$ or operators in $\mathcal{L}$, this support could be reduced to $\mathcal{R}_{\mathcal{L}\mathcal{S}_{O}}(X(v_{1})Z(u_{1}))$. Therefore, we have
$\mathcal{R}_{\mathcal{L}\mathcal{S}_{O}}(P)$ $\geq|B+S^{X}v_{1}+S^{Z}u_{1}|+|v_{m}+v_{L_{X}}|+|u_{m}+u_{L_{Z}}|+|v_{1}+v_{m}|+|u_{1}+u_{m}|$
$+\min(|v_{m}|,|\bar{X}|-|v_{m}|)+\min(|u_{m}|,|\bar{Z}|-|u_{m}|)+\mathcal{R}_{\mathcal{L}\mathcal{S}_{O}}(X(v_{1})Z(u_{1})).$ (91)
We do a case analysis based on $|v_{m}|,|u_{m}|$.
  1. 1.
    If $|v_{m}|\leq|\bar{X}|-|v_{m}|$ and $|u_{m}|\leq|\bar{Z}|-|u_{m}|$, we have
    $\mathcal{R}_{\mathcal{L}\mathcal{S}_{O}}(P)$ $\geq|B+S^{X}v_{1}+S^{Z}u_{1}|+|v_{1}+v_{m}|+|u_{1}+u_{m}|+|v_{m}|+|u_{m}|$ (92)
    $\geq|B+S^{X}v_{1}+S^{Z}u_{1}|+|v_{1}|+|u_{1}|$ (93)
    $\geq|B|\geq d.$ (94)
  2. 2.
    If $|v_{m}|\geq|\bar{X}|-|v_{m}|$ and $|u_{m}|\leq|\bar{Z}|-|u_{m}|$, we have
    $\mathcal{R}_{\mathcal{L}\mathcal{S}_{O}}(P)$ $\geq|u_{1}+u_{m}|+|u_{m}|+|\bar{X}|-|v_{m}|+\mathcal{R}_{\mathcal{L}\mathcal{S}_{O}}(v_{1}u_{1})$ (95)
    $\geq|u_{1}|+|\bar{X}|-|v_{1}|+\mathcal{R}_{\mathcal{L}\mathcal{S}_{O}}(v_{1}u_{1}).$ (96)
    Note that $\mathcal{R}_{\mathcal{L}\mathcal{S}_{O}}(v_{1}u_{1})+|u_{1}|\geq\mathcal{R}_{\mathcal{L}\mathcal{S}_{O}}(v_{1})$. Therefore $\mathcal{R}_{\mathcal{L}\mathcal{S}_{O}}(P)\geq|\bar{X}|-|v_{1}|+\mathcal{R}_{\mathcal{L}\mathcal{S}_{O}}(v_{1})\geq\mathcal{R}_{\mathcal{L}\mathcal{S}_{O}}(\bar{X})\geq d$.
  3. 3.
    The case $|v_{m}|\leq|\bar{X}|-|v_{m}|$ and $|u_{m}|\geq|\bar{Z}|-|u_{m}|$ follows similarly.
  4. 4.
    If $|v_{m}|\geq|\bar{X}|-|v_{m}|$ and $|u_{m}|\geq|\bar{Z}|-|u_{m}|$, we have
    $\mathcal{R}_{\mathcal{S}}(P^{\prime})$ $\geq|\bar{X}|-|v_{m}|+|\bar{Z}|-|u_{m}|+\mathcal{R}_{\mathcal{L}\mathcal{S}_{O}}(v_{1}u_{1})$ (97)
    $\geq|\bar{X}|-|v_{1}|+|\bar{Z}|-|u_{1}|+\mathcal{R}_{\mathcal{L}\mathcal{S}_{O}}(v_{1}u_{1})$ (98)
    $\geq\mathcal{R}_{\mathcal{L}\mathcal{S}_{O}}(\bar{X}\cdot\bar{Z})\geq d.$ (99)
Therefore we have $d(\bar{X}_{M}\mathcal{L}\mathcal{U})=\mathcal{R}_{\mathcal{L}\tilde{S}}(Z(B)q_{1}^{Z})\geq d$, as desired. ∎
A.3 Proofs for the Multi-block $XX$ Ancilla System
See 24
Proof.
Let $\mathcal{S}$ be the stabilizer group of the original codes and each ancilla qubit, including bridge qubits, in |$\ket{0}$, and $\mathcal{S}_{XX}$ be the stabilizer group of the joint merged code, Figure 6. Also let $\mathcal{L}$ be the group of logical operators of the unmeasured $12k-$ 1 qubit logical space and =$\mathcal{L}^{*}=\mathcal{L}\setminus\{I\}$. The versions of Lemmas 9 and 10 that apply to the joint code imply that the measurement fault distance is $\min(R,d_{Z}(\bar{Z}_{1}\mathcal{L}\mathcal{S}))$ and the logical fault distance is $\min(d_{Z}(\mathcal{L}^{*}\mathcal{S}),d_{X}(\mathcal{L}^{*}\mathcal{S}_{XX}))$. The original code $\mathcal{S}$ has code distance at least $d$ – if it did not, then the individual systems would not have fault distances at least $d$ – and so $d_{Z}(\bar{Z}_{1}\mathcal{L}\mathcal{S})\geq d$ and $d_{Z}(\mathcal{L}^{*}\mathcal{S})\geq d$. To complete the proof, it remains only to show that the $X$-distance of the joint merged code is good, i.e. $d_{X}(\mathcal{L}^{*}\mathcal{S}_{XX})\geq d$.
Since $|X(B)S_{1}|$ is at least the weight of factors in line 101, we proceed by ignoring the factors in line 102.